# Fine-Grained Cross-View Localization via Local Feature Matching and Monocular Depth Priors

### v1

@@ Abstract
We propose an accurate and highly interpretable fine-grained cross-view localization method that estimates the 3 Degrees of Freedom pose of a ground-level image by matching its local features with a reference aerial image. Previous methods typically transform the ground image into a bird’s-eye view (BEV) representation and then align it with the aerial image for localization. However, this transformation often leads to information loss due to perspective distortion or compression of height information, thereby degrading alignment quality with the aerial view. In contrast, our method directly establishes correspondences between ground and aerial images and lifts only the matched keypoints to BEV space using monocular depth prior. Notably, modern depth predictors can provide reliable metric depth when the test samples are similar to the training data. When the depth distribution differs, they still produce consistent relative depth, i.e., depth accurate up to an unknown scale. Our method supports both metric and relative depth. It employs a scale-aware Procrustes alignment to estimate the camera pose from the correspondences and optionally recover the scale when using relative depth. Experimental results demonstrate that, with only weak supervision on camera pose, our method learns accurate local feature correspondences and achieves superior localization performance under challenging conditions, such as cross-area generalization and unknown orientation. Moreover, our method is compatible with various relative depth models without requiring per-model finetuning. This flexibility, combined with strong localization performance, makes it well-suited for real-world deployment.
@@ 1 Introduction
Visual localization, a fundamental task in computer vision and mobile robotics, aims to estimate the camera pose with respect to a representation of the environment Thrun (2002). In dense urban areas, specialized positioning sensors, such as Global Navigation Satellite System (GNSS), have errors up to tens of meters Xia et al. (2021). Fine-grained cross-view localization, a recently emerging task, offers a promising complement. It estimates the 3 Degrees of Freedom (DoF) camera pose, i.e., 2D planar location and yaw orientation, by comparing its captured ground-level images to an aerial image of the surroundings, identified using the noisy position estimates from GNSS.
Previous approaches tackle the cross-view localization task with various formulations, such as matching global image descriptors Xia et al. (2023), sliding ground-view transformed bird’s-eye-view (BEV) features over aerial images Fervers et al. (2023b), and estimating homography Wang et al. (2024b). Significant effort has been devoted to how to aggregate information along the vertical (height) axis to transform the ground-view feature to BEV, such that the resulting BEV can align well with aerial imagery. Song et al. (2024) designs a learned module to refine the distortion caused by above-ground objects in homography. Wang et al. (2024a) uses attention to gather information from above-ground objects, while Xia and Alahi (2025) explicitly classifies features along height.
However, transforming the ground image to BEV still inevitably leads to information loss, as it may introduce ray-directional distortions on objects Song et al. (2024), or the model must discard or compress information along height Xia and Alahi (2025); Wang et al. (2024a). These limitations can degrade the semantic richness of the BEV feature, thereby impairing its alignment with aerial imagery. In contrast, we propose to directly establish correspondences between ground and aerial images, without explicitly selecting features along height. The correspondence between ground and aerial views is inherently many-to-one: rather than restricting the match of a rooftop in the aerial image to only the top of a building facade in the ground view, the entire facade can contribute to this association. On the one hand, the topmost point may not exhibit similar appearance in the aerial view. On the other hand, image matching has evolved beyond identifying visually similar content to capturing semantically consistent correspondences, as evidenced in visual grounding Li et al. (2019a); Yan et al. (2023); Yang et al. (2022) and cross-domain visual matching Ren et al. (2024).
Figure 1: Our method addresses fine-grained cross-view localization by matching local features between ground and aerial images. Using the depth from a monocular depth predictor, we lift only the matched ground keypoints to BEV space, and then jointly compute rotation, translation, and scale using these correspondences to align the BEV points with the aerial image. Our image matching is learned through weak supervision from the camera pose.
After matching image-plane correspondences, an intuitive approach to computing the 3-DoF camera pose is to propagate the correspondences into a shared BEV space. To directly bridge the ground view and BEV, previous methods rely on LiDAR Lee et al. (2025); Wang et al. (2023b) or monocular depth models Wang et al. (2025). Since LiDAR is not widely available, we opt for a vision-based solution. Recent monocular depth models Yang et al. (2024a, b); Piccinelli et al. (2025); Wang and Liu (2024) provide accurate metric depth when test scenes resemble their training distribution. When generalizing to a new distribution, absolute depth may be ambiguous, but these models still yield reliable non-metric relative depth, i.e., depth up to an unknown scale. This relative depth is still highly valuable, as it preserves the scene’s spatial structure. Notably, when humans self-localize using Google Satellite Maps Google (2024), they often explicitly match co-visible contents, e.g., road markings and buildings, across views, and then leverage the relative geometry of objects to find the ego-location, a process now replicated by our method. As shown in Fig. 1, our proposed method first identifies sparse correspondences between ground and aerial images, uses a monocular depth prior to map the matched ground keypoints into the BEV space, and employs scale-aware Procrustes alignment Umeyama (1991) to compute the rotation, translation and scale, between the BEV and aerial image.
@@
Concretely, our main contributions are: (1) We propose a simple and accurate fine-grained cross-view localization method that matches local features across views. Our method achieves superior localization accuracy under challenging conditions, such as cross-area generalization and unknown orientation. On the KITTI cross-area test set, it reduces the mean localization error by 26% compared to the previous state-of-the-art. (2) Our method establishes accurate correspondences between ground and aerial images. By analytically computing the pose from these correspondences, it delivers highly interpretable localization results. The feature matching is learned without explicit labels, relying only on weak supervision from the camera pose. (3) At inference time, our method is plug-and-play with different relative depth predictors, requiring no finetuning and resulting in only a minor drop in localization accuracy. Once reliable correspondences with the metric aerial imagery are established, our method can recover the scale of the relative depth.
@@ 2 Related Work
Fine-grained cross-view localization is a challenging task due to the drastically different viewpoints and the varying times at which ground and aerial images are captured. State-of-the-art approaches tackle this domain gap primarily by learning a deep network for camera pose estimation. Global descriptor-based methods Xia et al. (2022); Lentsch et al. (2023); Xia et al. (2023) leverage contrastive learning to pull the ground and aerial features to a common representation space, and compare the ground image descriptor to aerial descriptors extracted at different candidate poses. Geometry transformation-based methods Fervers et al. (2023b); Wang et al. (2023a); Sarlin et al. (2024); Wang et al. (2024a); Song et al. (2024); Shi et al. (2023); Wang et al. (2024b); Fervers et al. (2023a); Xia and Alahi (2025) first warp the ground image into a BEV representation, and then either perform a sliding-window search over the aerial image to find the best alignment Sarlin et al. (2024); Fervers et al. (2023b); Shi et al. (2023) or estimate the relative pose by directly matching features between the BEV and aerial view Song et al. (2024); Wang et al. (2024b, 2023a, a); Xia and Alahi (2025). Despite the increasing localization accuracy, most methods do not explicitly identify matched local features across views, offering limited interpretability. A recent work Xia and Alahi (2025) demonstrated, for the first time, the ability to find ground-aerial local feature correspondences by defining a dense 3D voxel space and querying image features for each voxel. However, this process is inefficient due to the sparse nature of 3D space, and this method performs poorly when camera orientation is unknown.
Ground-view visual localization typically follows a highly interpretable pipeline Sarlin et al. (2019); Sattler et al. (2016), which involves finding local image feature correspondences Sarlin et al. (2020); Sun et al. (2021) and then estimating camera pose using a geometry-based solver Lepetit et al. (2009). However, this pipeline becomes challenging in the cross-view setting due to the significant domain gap and the lack of labeled correspondences. We propose to learn such correspondences for cross-view localization without labels, facilitated by monocular depth priors.
Monocular depth estimation bridges 2D imagery and 3D scene. Early methods Li et al. (2019b); Fu et al. (2018); Liu et al. (2015); Roy and Todorovic (2016) demonstrated strong performance in predicting reliable depth on in-domain data. Later, unified loss functions and architectures Ranftl et al. (2021, 2022) enabled joint training on diverse datasets Geiger et al. (2013); Li and Snavely (2018); Nathan Silberman and Fergus (2012); Sturm et al. (2012), making generalizable relative depth prediction possible, i.e., estimating depth values that are accurate up to an unknown scale Wang and Liu (2024); Yang et al. (2024b). By finetuning on ground-truth depth, these methods can also produce accurate metric depth, i.e., depth values in absolute metric space Yang et al. (2024a, b); Bhat et al. (2023). Recently, training depth models with camera-awareness Zhu et al. (2024); Yin et al. (2023); Hu et al. (2024); Song et al. (2025); Piccinelli et al. (2024, 2025) has further advanced the feasibility of generalizable metric depth prediction. Considering all the variations of depth models, we design our cross-view localization method to support both metric and relative depth. The metric depth provides guidance on learning accurate local feature matching. Once trained, our model can operate in a plug-and-play manner at inference time with various depth predictors, including relative depth ones.
@@ 3 Methodology
Fine-grained cross-view localization aims to estimate the 3-DoF pose, i.e., the 2D metric location $\mathbf{t}\in\mathbb{R}^{2}$ and yaw orientation $o\in[-\pi,\pi)$, of a ground-level image $G$ by matching it to a geo-referenced aerial image $A$ covering the local surrounding area.
3.1 Motivation
Our goal is to solve fine-grained cross-view localization via local feature matching. Since existing datasets lack ground-truth correspondences between ground and aerial images, we learn the matching in a weakly supervised manner using only the pose. To support this process, we incorporate a pre-trained monocular depth model to predict a depth map of the ground image, i.e., $D=\mathcal{D}(G)$.
Importantly, monocular depth prediction is a geometrically ill-posed problem, as there is no unique mapping from a single image to absolute metric depth. State-of-the-art methods either learn generalizable metric depth prediction from diverse training data Piccinelli et al. (2025, 2024); Zhu et al. (2024), or predict depth that is accurate only up to a per-image unknown scale Wang and Liu (2024); Yang et al. (2024a) [^1] Recent work Yu et al. (2025) also explored modeling the relationship between relative and metric depth using a scale and a shift. As our goal is not metric depth prediction, we follow Eigen et al. (2014); Ding et al. (2024) and consider only the scale for simplicity. Therefore, our method is designed to use relative depth, while it also supports metric depth when available. Notably, different depth predictors can be seamlessly integrated into our method at inference time without the need for finetuning or retraining.
Figure 2: Our method first matches local features between ground and aerial images. The sampled correspondences are then lifted to the BEV space using monocular depth priors. By aligning these correspondences using scale-aware Procrustes alignment, we estimate the rotation, translation, and scale between the ground and aerial views.
As shown in Fig. 2, our method employs two feature extraction branches sharing the same architecture. Each branch includes a frozen pre-trained DINOv2 encoder Oquab et al. (2024) followed by a lightweight projection head, composed of several convolutional layers and a self-attention layer Vaswani et al. (2017). The branches map the ground image $G$ and aerial image $A$ to local image feature maps $\mathbf{F}_{G}$ and $\mathbf{F}_{A}$, respectively.
3.2 Pose Estimation
Given the ground and aerial feature maps $\mathbf{F}_{G}$ and $\mathbf{F}_{A}$, our goal is to estimate the 3-DoF metric camera pose by establishing correspondences between them. We first perform feature matching between $\mathbf{F}_{G}$ and $\mathbf{F}_{A}$, and subsequently leverage pre-trained monocular depth models to assign coordinates to the correspondences, which are then used to estimate the camera pose. Notably, reliable depth predictors are now available for both perspective images Wang and Liu (2024); Yang et al. (2024b) and panoramic images Wang and Liu (2024); Piccinelli et al. (2025).
Matching: We first compute the pairwise matching scores, $M$, between aerial and ground features $\mathbf{F}_{A}$ and $\mathbf{F}_{G}$ based on cosine similarity, $M=\operatorname{cosine}(\mathbf{F}_{A},\mathbf{F}_{G})/\tau$, where $\tau$ is a temperature parameter. Following DeTone et al. (2018); Sarlin et al. (2020), we append a learnable dustbin to both the rows and columns of $M$, allowing the model to reject uncertain or unmatched keypoints. On this extended matching score matrix $M^{*}$, we apply dual softmax normalization over rows and columns to obtain a match probability matrix $\hat{M}^{*}$,
@@ 3.2 Pose Estimation
Finally, we drop the dustbin row and column from $\hat{M^{*}}$, and sample $N$ correspondences for pose estimation. We denote the matching probability of the sampled correspondences as $w_{n}$, $n\in[1,...,N]$, which will be used as the weight in the Procrustes alignment.
Coordinate assignment: Given the $N$ sampled correspondences, we denote the metric coordinates of the 2D point associated with the $n$-th aerial feature as $(x_{n}^{A},y_{n}^{A})$, with the origin defined at the center of the aerial image. For the corresponding ground feature, we retrieve its depth from $D$ and use its associated ray direction to compute its 3D location in space. The resulting 3D point is denoted as $(x_{n}^{G},y_{n}^{G},z_{n}^{G})/s$, with the origin defined at the ground camera. Here, $s$ is an unknown scale factor from the ground coordinate system to the aerial metric space, and $s=1$ when a reliable metric depth model is used. As motivated in Sec. 1, we retain all matched ground points without explicitly selecting them based on their height (z-coordinate). Our experiments will show that explicitly selecting the topmost point does not improve performance.
Scale-aware Procrustes alignment: As shown in Fig. 2 right, once we have the point correspondences $\{(x_{n}^{A},y_{n}^{A}),(x_{n}^{G},y_{n}^{G},z_{n}^{G})/s\}$, and their matching probabilities $w_{n}$, the rotation $\mathbf{R}$, translation $\mathbf{t}$, and the scale $s$ between the ground points and the metric space aerial points can be computed analytically in a differentiable manner using 2D scale-aware Procrustes alignment Umeyama (1991)  [^2] In Xia and Alahi (2025); Barroso-Laguna et al. (2024), all coordinates are in metric space, so they use orthogonal Procrustes Gower (1975) and ignore scale..
@@
In Eq. 2, the better matched points will contribute more to the centroids of the point sets. Then, we center the point sets and compute their covariance matrix,
$\tilde{\mathbf{Q}}=\mathbf{Q}-\bar{\mathbf{Q}},\quad\tilde{\mathbf{P}}=\mathbf{P}-\bar{\mathbf{P}},\quad\mathbf{C}=\sum_{n=1}^{N}w_{n}\left(\tilde{\mathbf{P}}_{n}\right)\left(\tilde{\mathbf{Q}}_{n}\right)^{\top},$ (3)
where $\tilde{\mathbf{P}}_{n}$ and $\tilde{\mathbf{Q}}_{n}$ are the $n$-th points in the centered point sets $\tilde{\mathbf{Q}}$ and $\tilde{\mathbf{P}}$. The covariance matrix $\mathbf{C}$ captures how the centered ground points relate to the centered aerial points, encoding the direction and strength of their mutual variation.
Next, we perform Singular Value Decomposition (SVD) on $\mathbf{C}$, $\mathbf{C}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$, and the optimal rotation can be recovered as $\mathbf{R}=\mathbf{V}\mathbf{U}^{\top}$, which derives the estimated yaw orientation $o$. The scale and translation from the ground points $\mathbf{P}$ to the aerial points $\mathbf{Q}$ is then computed as,
@@
Recall that $\mathbf{P}=\mathbf{P}^{*}/s$. Substituting this into Eq. 4, the estimated scale $s^{*}$ can be rewritten in terms of the (unscaled) metric coordinates $\mathbf{P}^{*}$, with $\mathbf{\Sigma}^{*}$ denoting the singular value matrix computed from the SVD between $\mathbf{Q}$ and $\mathbf{P}^{*}$:
$s^{*}=\frac{\operatorname{Tr}(\mathbf{\Sigma}^{*}/s)}{\sum_{n=1}^{N}w_{n}\left\|\tilde{\mathbf{P}^{*}}^{n}/s\right\|^{2}}=\frac{\operatorname{Tr}(\mathbf{\Sigma}^{*})/s}{\sum_{n=1}^{N}w_{n}\left\|\tilde{\mathbf{P}^{*}}^{n}\right\|^{2}/s^{2}}=\frac{s\operatorname{Tr}(\mathbf{\Sigma}^{*})}{\sum_{n=1}^{N}w_{n}\left\|\tilde{\mathbf{P}^{*}}^{n}\right\|^{2}}.$ (5)
@@
Assumptions: In practice, the implication holds only when the transformation between aerial and ground points is an exact similarity transform. On the one hand, it requires that the depth prior and scene geometry are well-aligned, i.e., the relative depth estimated from the ground image must correspond to the aerial layout, and the scene geometry must be sufficiently structured to support stable alignment. On the other hand, the estimated correspondences must be highly accurate. Assuming that relative depth predictions are generally reliable, we focus on learning accurate correspondences. Our experiments demonstrate that, once accurate local feature correspondences are established, our model can consistently estimate pose, even when different scales are applied to the ground points.

### v2

# Loc2: Interpretable Cross-View Localization via Depth-Lifted Local Feature Matching

@@ Abstract
We propose an accurate and interpretable fine-grained cross-view localization method that estimates the 3 Degrees of Freedom (DoF) pose of a ground-level image by matching its local features with a reference aerial image. Unlike prior approaches that rely on global descriptors or bird’s-eye-view (BEV) transformations, our method directly learns ground–aerial image-plane correspondences using weak supervision from camera poses. The matched ground points are lifted into BEV space with monocular depth predictions, and scale-aware Procrustes alignment is then applied to estimate camera rotation, translation, and optionally the scale between relative depth and the aerial metric space. This formulation is lightweight, end-to-end trainable, and requires no pixel-level annotations. Experiments show state-of-the-art accuracy in challenging scenarios such as cross-area testing and unknown orientation. Furthermore, our method offers strong interpretability: correspondence quality directly reflects localization accuracy and enables outlier rejection via RANSAC, while overlaying the re-scaled ground layout on the aerial image provides an intuitive visual cue of localization accuracy.
@@ 1 Introduction
Visual localization, a fundamental task in computer vision and mobile robotics, aims to estimate the camera pose with respect to a representation of the environment (Thrun, 2002). In dense urban areas, specialized positioning sensors, such as Global Navigation Satellite System (GNSS), have errors up to tens of meters (Xia et al., 2021). Fine-grained cross-view localization has recently emerged as a promising complement. It estimates the 3 Degrees of Freedom (DoF) camera pose, i.e., 2D planar location and yaw orientation, by comparing captured ground-level images to an aerial image of the surroundings, identified using coarse GNSS measurements.
The key to cross-view localization is to establish associations between ground-level and aerial images. However, the extreme visual differences between the two views make it challenging for recent advances in image matching (Leroy et al., 2024; Sun et al., 2021; Wang et al., 2025) to produce reliable correspondences, and there is no ground-aerial pixel-level ground truth to finetune these methods. In practice, the majority of cross-view localization approaches seek global alignment between two views by matching global image descriptors (Xia et al., 2023) or aligning ground-view transformed bird’s-eye-view (BEV) features with aerial images (Fervers et al., 2023b; Wang et al., 2024b). However, these methods offer limited interpretability, as they cannot explicitly identify which objects in the ground view correspond to those in the aerial view.
Figure 1: Loc2: Interpretable cross-view localization via local feature matching. Loc2 establishes accurate correspondences between aerial and ground views, with colors indicating distinct correspondence regions. Using the estimated rotation, translation, and scale, the ground view is warped onto the aerial image, providing a visual interpretation of localization quality.
Recently, FG2 (Xia & Alahi, 2025) demonstrated the feasibility of establishing local feature correspondences between ground and aerial images by propagating BEV point matches to the image plane. However, performing the matching in BEV is suboptimal, since warping a ground image into BEV inevitably introduces ray-directional distortions (Song et al., 2024) and discards information along the height dimension (Xia & Alahi, 2025; Wang et al., 2024a). This loss of information hinders matching to aerial images and, as a result, degrades localization performance, especially in challenging settings such as unknown camera orientation.
Therefore, we propose to establish correspondences directly between ground and aerial images. Instead of relying on pixel-level annotations (Leroy et al., 2024; Wang et al., 2025), our approach is weakly supervised using only 3-DoF camera poses. To infer the pose from image-plane correspondences, we lift the sampled ground points into the BEV space with the help of monocular depth models (Yang et al., 2024a; b; Piccinelli et al., 2025; Wang & Liu, 2024). In unseen environments, usually only the non-metric relative depth prediction, i.e., depth up to an unknown scale, is reliable. Therefore, our method supports both metric and relative depth, and jointly estimates the camera pose and depth scale using scale-aware Procrustes alignment (Umeyama, 1991).
As shown in Fig. 1, our method offers superior interpretability. Since the pose is computed analytically from the matches, the quality of local feature correspondences directly reflects localization accuracy, and we find that the number of inlier correspondences shows a strong correlation with localization accuracy. Moreover, by overlaying the scaled, rotated, and translated ground BEV layout onto aerial images, our method provides an additional visual cue of localization accuracy.
@@
Concretely, our main contributions are: (1) We propose a simple and accurate fine-grained cross-view localization method that matches local features across views. Our method achieves superior localization accuracy under challenging conditions such as cross-area generalization and unknown orientation. (2) Our method is highly interpretable. The pose is computed analytically from the estimated correspondences, which also enable outlier filtering via RANSAC. When relative depth is used at inference, our method estimates its scale, and the alignment between the re-scaled ground layout and the aerial image further provides a visual cue of localization quality. (3) By leveraging differentiable scale-aware Procrustes alignment, our local feature matching becomes end-to-end trainable using only camera pose supervision, without requiring pixel-level annotations.
@@ 2 Related Work
Fine-grained cross-view localization is challenging due to the drastic viewpoint differences and asynchronous capture times between ground and aerial imagery. State-of-the-art approaches tackle this domain gap primarily by learning a deep network for camera pose estimation. Global descriptor-based methods (Xia et al., 2022; Lentsch et al., 2023; Xia et al., 2023) leverage contrastive learning to pull the ground and aerial features to a common representation space, and compare the ground image descriptor to aerial descriptors extracted at different candidate poses. Geometry transformation-based methods (Fervers et al., 2023b; Wang et al., 2023; Sarlin et al., 2024; Wang et al., 2024a; Song et al., 2024; Shi et al., 2023; Wang et al., 2024b; Fervers et al., 2023a; Xia & Alahi, 2025) first warp the ground image into a BEV representation, and then either perform a sliding-window search over the aerial image to find the best alignment (Sarlin et al., 2024; Fervers et al., 2023b; Shi et al., 2023) or estimate the relative pose by matching features between the BEV and aerial view (Song et al., 2024; Wang et al., 2024b; 2023; a; Xia & Alahi, 2025). Despite the increasing localization accuracy, most methods do not explicitly identify matched local features across views, offering limited interpretability. A recent work (Xia & Alahi, 2025) demonstrated, for the first time, the ability to find ground-aerial local feature correspondences by defining a dense 3D point cloud and querying image features for each point. However, this process is inefficient due to the sparse nature of 3D space, and this method performs poorly when camera orientation is unknown.
Ground-view visual localization typically follows a highly interpretable pipeline (Sarlin et al., 2019; Sattler et al., 2016), which involves finding local feature correspondences (Sarlin et al., 2020; Sun et al., 2021) and then estimating camera pose using a geometry-based solver (Lepetit et al., 2009). However, this pipeline becomes challenging in the cross-view setting due to the significant domain gap and the lack of labeled correspondences. We propose to learn such correspondences for cross-view localization using only camera pose supervision, facilitated by monocular depth priors.
@@ 3 Methodology
Fine-grained cross-view localization estimates the 3-DoF pose, i.e., the 2D metric location $\mathbf{t}\in\mathbb{R}^{2}$ and the yaw orientation $o\in[-\pi,\pi)$, of a ground-level image $G$ by matching it to a geo-referenced aerial image $A$ covering the local area. As shown in Fig. 2, our method first establishes correspondences between $G$ and $A$ (Sec. 3.1), then lifts the matched ground points to the BEV using the predicted depth map $D=\mathcal{D}(G)$ from off-the-shelf monocular depth models, and finally computes the 3-DoF pose via scale-aware Procrustes alignment (Sec. 3.2). This fully differentiable pipeline enables end-to-end learning of image-plane matching under weak supervision from camera poses.
Figure 2: Overview of our proposed method. Our method first matches local features between ground and aerial images. The matched ground points are then lifted to the BEV space using monocular depth priors. By aligning these correspondences using scale-aware Procrustes alignment, we estimate the rotation, translation, and scale between the ground and aerial views.
3.1 Local Feature Matching
We estimate the 3-DoF metric camera pose by learning image-plane local feature correspondences between ground and aerial images. Our model has two feature extraction branches that share the same architecture. Each branch consists of a frozen DINOv2 feature extractor (Oquab et al., 2024) followed by our lightweight projection head, which is composed of several convolutional layers and a self-attention layer (Vaswani et al., 2017). The two branches map the aerial image $A$ and the ground image $G$ to feature maps $\mathbf{F}_{A}$ and $\mathbf{F}_{G}$, respectively.
Afterwards, we compute the pairwise matching scores $M$ between $\mathbf{F}_{A}$ and $\mathbf{F}_{G}$ using cosine similarity, $M=\operatorname{cosine}(\mathbf{F}_{A},\mathbf{F}_{G})/\tau$, where $\tau$ is a temperature parameter. Following DeTone et al. (2018); Sarlin et al. (2020), we append a learnable dustbin to both the rows and columns of $M$, allowing the model to reject uncertain or unmatched points. On this extended matching score matrix $M^{*}$, we apply dual softmax normalization over rows and columns to obtain a match probability matrix $\hat{M}^{*}$,
@@ 3.1 Local Feature Matching
Finally, we drop the dustbin row and column from $\hat{M^{*}}$, and sample $N$ correspondences for subsequent pose estimation. We denote the matching probability of the sampled correspondences as $w_{n}$, $n\in[1,...,N]$, which will be used as the weight in the scale-aware Procrustes alignment.
3.2 Camera Pose Estimation
Coordinate assignment: Given the $N$ sampled correspondences, we leverage off-the-shelf monocular depth estimators to assign coordinates to these correspondences, which are then used to estimate the camera pose. Importantly, monocular depth prediction is a geometrically ill-posed problem, as there is no unique mapping from a single image to absolute metric depth. State-of-the-art methods either learn generalizable metric depth from diverse training data (Piccinelli et al., 2025; Zhu et al., 2024), or predict depth that is accurate only up to an unknown per-image scale (Wang & Liu, 2024; Yang et al., 2024a). To generalize across diverse scenarios and image types, our method accommodates both metric and relative depth, and explicitly estimates a scale factor $s$ to convert relative depth into metric space  [^1] Recent work (Yu et al., 2025) models relative depth using a scale and a shift. As our goal is not metric depth prediction, we follow Eigen et al. (2014); Ding et al. (2024) and consider only the scale for simplicity.. This scale estimation leverages the metric information, i.e., meters per pixel, available in geo-referenced aerial imagery.
Specically, we denote the metric coordinates of the 2D point associated with the $n$-th aerial feature as $(x_{n}^{A},y_{n}^{A})$, with the origin defined at the center of the aerial image. For the corresponding ground feature, we retrieve its depth from $D$ and use its associated ray direction to compute its 3D location in space. The resulting 3D point is denoted as $(x_{n}^{G},y_{n}^{G},z_{n}^{G})/s$, with the origin defined at the ground camera. Here, $s$ is an unknown scale factor from the ground coordinate system to the aerial metric space, and $s=1$ when a reliable metric depth model is used. We retain all matched ground points without explicitly selecting them based on their height (z-coordinate). Our ablation study will show that explicitly selecting the topmost point does not improve performance.
Scale-aware Procrustes alignment: As shown in Fig. 2 right, once we have the point correspondences $\{(x_{n}^{A},y_{n}^{A}),(x_{n}^{G},y_{n}^{G},z_{n}^{G})/s\}$, and their matching probabilities $w_{n}$, we can compute the rotation $\mathbf{R}$, translation $\mathbf{t}$, and the scale $s$ between the ground points and the metric space aerial points analytically in a differentiable manner using 2D scale-aware Procrustes alignment (Umeyama, 1991)  [^2] In Xia & Alahi (2025); Barroso-Laguna et al. (2024), all coordinates are in metric space, so they use orthogonal Procrustes (Gower, 1975) and ignore scale..
@@ 3.2 Camera Pose Estimation
In Eq. 2, the better matched points will contribute more to the centroids of the point sets. Then, we center the point sets and compute their covariance matrix $\mathbf{C}$,
$\mathbf{C}=\sum_{n=1}^{N}w_{n}\left(\tilde{\mathbf{P}}_{n}\right)\left(\tilde{\mathbf{Q}}_{n}\right)^{\top},$ (3)
where $\tilde{\mathbf{P}}_{n}$ and $\tilde{\mathbf{Q}}_{n}$ are the $n$-th points in the centered point sets $\tilde{\mathbf{Q}}=\mathbf{Q}-\bar{\mathbf{Q}}$ and $\tilde{\mathbf{P}}=\mathbf{P}-\bar{\mathbf{P}}$. The covariance matrix $\mathbf{C}$ captures how the centered ground points relate to the centered aerial points, encoding the direction and strength of their mutual variation.
Next, we perform Singular Value Decomposition (SVD) on $\mathbf{C}$, where $\mathbf{C}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$, and the optimal rotation can be recovered as $\mathbf{R}=\mathbf{V}\mathbf{U}^{\top}$, which derives the estimated yaw orientation $o$. The scale and translation from the ground points $\mathbf{P}$ to the aerial points $\mathbf{Q}$ is then computed as,
@@
Recall that $\mathbf{P}=\mathbf{P}^{*}/s$. Substituting this into Eq. 4, the estimated scale $s^{*}$ can be expressed using the (unscaled) metric coordinates $\mathbf{P}^{*}$, with $\mathbf{\Sigma}^{*}$ being the singular value matrix computed from the SVD between $\mathbf{Q}$ and $\mathbf{P}^{*}$:
$s^{*}=\frac{\operatorname{Tr}(\mathbf{\Sigma}^{*}/s)}{\sum_{n=1}^{N}w_{n}\left\|\tilde{\mathbf{P}^{*}}^{n}/s\right\|^{2}}=\frac{s\operatorname{Tr}(\mathbf{\Sigma}^{*})}{\sum_{n=1}^{N}w_{n}\left\|\tilde{\mathbf{P}^{*}}^{n}\right\|^{2}}.$ (5)